OpenStudy (anonymous):

what is the value of the expression 2x+15-3x when x=4?

OpenStudy (stormfire1):

Well, 2(4) + 15 - 3(4) = ?

OpenStudy (unklerhaukus):

\[8+15-12= ?\]

OpenStudy (stormfire1):

23 - 12 = ?